OpenStudy (anonymous):

Which ordered pair is a solution of the equation 2x − y = 9?

OpenStudy (anonymous):

do u hv other info?

OpenStudy (anonymous):

just substitute the pairs and the will give 9 will be the pair

OpenStudy (anonymous):

@kpowell032

OpenStudy (anonymous):

for example: the first pair(-4,1) 2*-4-1=9 -9 does not =9 so its not the pair do the same for the other pairs